OpenStudy (anonymous):

The substitution method to find all solutions of the system of equations to solve x^2 + y= 9 and. x-y+3=0. The answer are in brackets. I need help!!!! :(

OpenStudy (anonymous):

You can use equation (2) to write y=x+3. Then, you can substitute this into eqn (1) and solve.

OpenStudy (anonymous):

And the bfsgd. What do I do after that?

OpenStudy (anonymous):

If you do that you get \[x^2 + x - 6 = 0\] which can be factorised.

OpenStudy (anonymous):

bfsgd, how do you solve the whole problem? I really don't get it.

OpenStudy (anonymous):

So then \[x^2 + x - 6 = (x+3)(x-2) = 0\] so x=-3 or x=2. Then we can find out the y coordinates too. For x=-3, y=0 and for x=2, y=5.